Grade 12th passPhysical Chemistry

6.9 g of metal carbonate were dissloved in 60 mL of 2 N HCl. The excess acid was neutralized by 20 mL of 1 N NaOH. What is the equivalent wt. of metal.

Tarun Sharma
10 Years agoGrade 12th pass

1 Answer

Vikas TU
10 Years ago
let the metal carbonate be:
M2(CO3)X
molar mass = 2M + 60x
moles = 6.9/( 2M + 60x)
 
Now,
Molarity = Normality/V.F
            = 2/1 = 2
MV = moles
moles = (2*60-20*1)/1000 = 100/1000 = 0.1 { 20 mm neutralized by OH-}
 6.9/Z  =  0.1
Z = 6.9/0.1
Z  = 69
Molar mass of metal carbonate is  hence, 69.
moles of metal = 0.1*2 = 0.2
Eq. Weight = moles x valence factor.
                = 0.2 x 2
                = 0.4